OpenStudy (anonymous):

solve 5+4x+2 is greater than or equal to x+7

OpenStudy (anonymous):

5+4x+2 >= x+7 Simplify to 7+4x >= x+7 Simplifies to 3x>=0 x>=0 Strange.

OpenStudy (anonymous):

thanks